OpenStudy (anonymous):

Choose the equation of the vertical line that passes through the point (-2, -1).

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-2, -1).

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

y=-1 is a horizontal line.

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3.

OpenStudy (anonymous):

y-4=3(x-2) y=3x-2

OpenStudy (anonymous):

do you do flvs work?

OpenStudy (anonymous):

flvs?

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6.

OpenStudy (anonymous):

perpendicular so slope is m=-1/3 y-2=(-1/3)(x-6) y=(-1/3)x+4

OpenStudy (anonymous):

Consider the equation 3x + 5y = 45.

OpenStudy (anonymous):

Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

hello